OpenStudy (vshiroky):

Simplify and make your own your own 56x^7+21x^4+63x^3

OpenStudy (alekos):

You can start off by dividing out the x^4 as a factor

OpenStudy (littlebird):

Greatest common factor is actually 7x^3

OpenStudy (alekos):

I meant x^3

OpenStudy (jhannybean):

Yes, that is right. 56, 21, and 63 are all multiples of \(\color{red}{7}\). \(x^3) = \color{red}{x^3} \cdot 1\) \(x^7 = (\color{red}{x^3})(x^4)\) \(x^4 = (\color{red}{x^3})(x^1)\)

OpenStudy (vshiroky):

\[7x^3(8x^4+3x+9) \] Can anyone confirm this?

OpenStudy (vshiroky):

@Jhannybean

OpenStudy (vshiroky):

@ganeshie8

OpenStudy (jhannybean):

Good job @vshiroky :)